You (US company) imported Earth Moving Equipment (EME) from Australia. You imported EME at US$ 300 (with Cash) on Dec 1, 2018. On Dec 15, 2018, you sold EME to Australia at A$400 (Australian $) in AR. The exchange rate on Dec 15, 2018 was 2 A$/US$. The exchange rate on Dec 31, 2018 was 4 A$/US$. On Feb 1, 2019, your customer paid you in full. The exchange rate on Feb 1 was 1 A$/US$. What were NI in 2018 and 2019, respectively? 100, 200 -100, 300 -200, 200 100, 100
Answer #1

Values given in the Question:

Purchase Price of EME as on 1st Dec, 2018 by US Co. = US $ 300

Sale Price of EME as on 15th Dec, 2018 in A$ = A$ 400

Date of Payment= 1st Feb, 2019

The exchange rate on:

Dec 15, 2018 = 2 A$/US$

Dec 31, 2018 = 4 A$/US$

Feb 1 = 1 A$/US$.

By using the values given above:

NI/ NL in 2018:

Sale Value in US $ on 15th Dec, 2018 = A$400/2 = US $ 200

Les Purchase Price = US $ 300

Net Loss on sale (A) = US $ 100

Value of Amount receivables as on 15th Dec, 2018 = A$ 400(Sale Price)

Value as per exchange rate as on 31st Dec, 2018 = A$ 400= US$ 400/4 = US$ 100

Amount Receivable on 15th Dec, 2018 = US $ 200

Amount Receivable on 31st Dec, 2018 = US $ 100

Foreign Exchange Loss (unrealised) on 31st Dec, 2018 (B) = US $ 100

Net Loss in 2018(A)+(B) = US $ 200 i.e. NI is -US $ 200

NI/ NL in 2019:

Amount Receivable in US $ as on 31st Dec, 2018 = US $ 100

Amount actually realised in US $ on 1st Feb, 2019 = US $ 400 (when exchange rate was 1A$/1US$)

Therefore, NI in 2019 = US $ 300